OpenStudy (anonymous):

simplify 0.3x^2+0.6x^2-0.5x^2

Directrix (directrix):

.3x^2 + 0.6x^2 - 0.5x^2 --------------------- Factor out the common x^2 term to get: x^2( .3 + .6 - .5)

Directrix (directrix):

@Jasmineplayhouse What is this sum: .3 + .6 - .5 = ?

OpenStudy (anonymous):

4

Directrix (directrix):

The integer 4 or the decimal .4 (read as: four tenths)?

OpenStudy (anonymous):

.4

Directrix (directrix):

Great. Now, let us put the problem back together.

Directrix (directrix):

.3x^2 + 0.6x^2 - 0.5x^2 = x^2( .3 + .6 - .5) = x^2 (.4) = .4 x^2 --->

OpenStudy (anonymous):

Got it! thanks

Directrix (directrix):

Happy to help.